# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
where $\beta$ describes substrate friction and $\eta$ describes viscous dissipation by beam bending. Non-reciprocity $\zeta$ introduces a complex component to $\lambda(q)$, causing advection even in highly viscous environments. The dispersion relation Eq. (2) now provides a suggestive scaling argument for the onset of snapping under confinement.
When a beam of length $L$ is confined, Euler buckling can be predicted by balancing line tension against bending elasticity to give $\Lambda/A\sim 1/L^{2}$. Equation (2) suggests that the onset of snapping from within the buckled state comes from a second balance, namely between restoring elasticity and non-reciprocity: $\zeta/A\sim L$. To connect this continuum expression to our experimental parameters, we note that $\zeta$ is analogous to the bending modulus $A$ but coupled to a single extra gradient. The microscopic passive stiffness $B$ coarse-grains to the continuum $A$ as $A\sim Bl_{0}$, with $l_{0}$ the length of a single linkage. By analogy, $k^{\mathrm{o}}$—which responds to microscopic angular gradients—coarse-grains to $\zeta$ as $\zeta\sim k^{o}l_{0}^{2}$ [SM §IV]. Expressing this second balance with our microscopic parameters, we hypothesize that the onset of snapping occurs when $k^{o}/B\sim 1$.
Motivated by the observation that snapping is dominated by the lowest modes of the beam, to capture the essence of the transition we therefore construct a chain with only two angular degrees of freedom—an odd von Mises truss [Fig. 3(a), Supplementary Video 3] [34]. Compressing the truss and increasing $k^{\mathrm{o}}$, we find the same polarization and transition to self-snapping observed in our longer chains [Fig. 3(b)]. Tracking the snapping frequency $\omega$ as $k^{\mathrm{o}}$ is increased, we can indeeed collapse data for multiple passive stiffnesses $B$ onto a single critical threshold $k^{o}_{c}$ in $k^{o}/B$ [Fig. 3(c)]. However, we also find that oscillations do not emerge at a finite frequency, as per a standard Hopf bifurcation [29]. Instead, their frequency grows as $\omega\sim\sqrt{k^{\mathrm{o}}-k^{\mathrm{o}}_{\mathrm{c}}}$, i.e. a square-root singularity that is the hallmark of an exceptional transition [35].
Figure 3: The odd von Mises truss is a minimal model of non-reciprocal buckling. (a, b) We capture non-reciprocal buckling in an odd von Mises truss with clamped ends. We compress this truss a distance $\text{d}l$ from its rest length. As $k^{\mathrm{o}}$ increases, the truss polarizes, and then begins to self-snap. (a) The internal angles of the truss ($\theta_{\mathrm{L}},\theta_{\mathrm{LC}},\theta_{\mathrm{RC}},\theta_{\mathrm{R}}$) with experimental realization below. (b) The non-reciprocal torques at each vertex during a single snap-through. (c) We measure the self-snapping frequency $\omega$ as a function of $k^{\mathrm{o}}$ for a range of passive bending stiffness $B$: $B=8.5\text{ mNm}$ and $B=6.8\text{ mNm}$ (circles) are implemented via electronic feedback, with $B=20\text{ mNm}$ (squares) implemented via a silicone elastomer. Rescaling the activity by this bending rigidity, $k^{\mathrm{o}}/B$, and the snapping frequency $\omega$ by the characteristic dissipative timescale $\omega\Gamma/B$ collapses these data onto a single curve given by Eq. 4. We find a square-root growth of frequencies $\omega\sim\sqrt{k^{\mathrm{o}}-k^{\mathrm{o}}_{c}}$ beyond a critical activity $k^{\mathrm{o}}_{c}$ that is characteristic of an exceptional transition. (d, e) Our theory Eq. 3 gives a complete phase diagram of the bifurcations in this minimal system. This shows a stable region (gray) and buckling phases: at lower compression a simple buckling (blue), then at higher compressions buckling with transient jackknifing (green), along with the snapping phase where the non-reciprocity creates a limit cycle in phase space (orange). Crucially these four regions encircle a critical exceptional point, CEP, red dot in panel (d). The purple and brown dots correspond to a cusp bifurcation and a global bifurcation where the narrow region of multi-stable buckling and snapping collapses to a global SNIC bifurcation.
To explore the origin of this exceptional transition, we develop a detailed model of the truss dynamics. Each node in the truss experiences a non-reciprocal angular bond tension $\tau_{i}=k^{\mathrm{o}}(\delta\theta_{i+1}-\delta\theta_{i-1})$, passive bending elasticity $\tau_{i}=-B\delta\theta_{i}$, and viscous dissipation $\tau_{i}=-\Gamma\delta\dot{\theta}_{i}$. We additionally allow for a finite longitudinal spring stiffness $k$, which is typically large, $k\gg Bl_{0}^{2}$ for a linkage spacing $l_{0}$. By using the principle of virtual work, we derive equations of motion (SM §III, IV), which are conveniently expressed in terms of the symmetric $\theta_{\text{S}}=\frac{1}{2}(\theta_{\text{L}}+\theta_{\text{R}})\text{,}$ and anti-symmetric $\theta_{\text{A}}=(\theta_{\text{L}}-\theta_{\text{R}})$ combinations of the end node angles $\theta_{L,R}$:
@@
where $\bm{\theta}=(\theta_{S},\theta_{A})$ and the coefficients ${\bf A}_{m}$ are functions of $k$, $k^{\mathrm{o}}$ and the compression $\mathrm{d}l$. The up/down symmetry of buckling imposes a $\mathbb{Z}_{2}$ equivariance $(\theta_{S},\theta_{A})\rightarrow(-\theta_{S},-\theta_{A})$ on the dynamics Eq. (3), forcing odd powers of $\theta$ only, in analogy to a classic pitchfork bifurcation. The additional symmetry under simultaneous reflection and a reversal in the sense of non-reciprocity, $(\theta_{S},\theta_{A},k^{\mathrm{o}})\rightarrow(\theta_{S},-\theta_{A},-k^{\mathrm{o}})$, constrains the functional forms of ${\bf A}_{m}$ and forces $k^{\mathrm{o}}$ to appear only in the off-diagonal elements of the linearized response matrix $J$ (SM §III).
The bifurcation structure of Eq. (3) is shown in Fig. 3(d). We find a line of exceptional transitions, which meet a second line of pitchfork/Hopf bifurcations at a single point $(k_{\mathrm{c}}^{\mathrm{o}},\mathrm{d}l_{\mathrm{c}})=({1}/{2\sqrt{5}}+{108}/{\sqrt{5}k},{45}/{2k})$ in the limit of large longitudinal stiffness $k$. At this point we have a double-zero eigenvalue in $J$: the combination of the critical phenomena of buckling and the non-Hermitian physics of non-reciprocity meet at a critical exceptional point (CEP). Such CEPs have recently been proposed to lead to run-and-chase dynamics in mean-field models of non-reciprocal spin [24, 33, 36] and reaction-diffusion [23, 21, 22] systems. Here, we show an analogous behavior in a non-reciprocal beam. Expanding $J$ about the CEP as $k^{\mathrm{o}}=k^{\mathrm{o}}_{\mathrm{c}}+\delta k^{\mathrm{o}}$ and $\mathrm{d}l=\mathrm{d}l_{\mathrm{c}}+\delta l/k$ and taking the rigid linkage limit $k\to\infty$ we find
which has eigenvalues $\lambda\sim\delta l+\sqrt{\delta l+\delta k^{\mathrm{o}}}$, in excellent agreement with our experimentally obtained snapping frequencies [Fig. 3(c)].
Our experimental cycle—first compression, next increasing activity, and finally relaxation—traces a loop that encircles the CEP in Fig. 3(d). Along this loop Eq. (3) predicts four phases, labeled (I)—(IV) in Fig. 3(d,e), which collectively form a $\mathbb{Z}_{2}$-symmetric Bogdanov-Takens bifurcation [37]. Here, the $\mathbb{Z}_{2}$ symmetry stems from the physics of buckling, but such discrete symmetries are prevalent in many minimal condensed matter systems [36]. (I) labels the flat unbuckled state [Fig 3(d, e) gray region]. Starting from the flat state $(k^{\mathrm{o}},\mathrm{d}l)=(0,0)$, upon increasing $\mathrm{d}l$ two symmetrically placed stable fixed points emerge as $\theta_{S}\sim\pm\sqrt{\delta l}$ in a pitchfork bifurcation, leaving the flat state as a marginally stable saddle point. Initially $\theta_{A}=0$ because the buckled state is left-right symmetric. As non-reciprocity $k^{\mathrm{o}}$ increases, the truss polarizes, and these stable points acquire nonzero $\theta_{A}$. The region with two stable buckled states is labeled (II) [Fig. 3(d,e) blue]. For sufficiently large compressions $\mathrm{d}l$ a new pair of saddle fixed points emerge from the flat state, representing transient jacknifed truss configurations. The flat state is now fully unstable. This region we denote (III), see Fig. 3(d,e), green. Upon further increase of $k^{\mathrm{o}}$ we cross the exceptional transition. At the nonlinear level this transition to snapping manifests as a global saddle-node invariant circle (SNIC) bifurcation, in which each pair of stable/unstable fixed points merge to leave behind a single limit cycle. This snapping phase we denote (IV), see Fig. 3(d,e), orange. Finally, as compression $\mathrm{d}l$ is decreased this limit cycle vanishes in a supercritical Hopf bifurcation.
Our theory Eq. (3) predicts that close to the CEP, the SNIC bifurcation is in fact composed of several more intricate bifurcations, including a region of dynamical multistability, in which snapping and fixed buckled states exist simultaneously. In SM §III we characterize this region of global bifurcations precisely. For the truss such multistability exists in a narrow region of parameter space; the purple and brown dots in Fig. 3(c) denote the cusp bifurcation where region (II) disappears and the point where the global multistable region collapses into a SNIC respectively. This region of multistability is sufficiently small that it is not visible on this plot. Overall the linearized dynamics of Eq. (4) provides an excellent description of our experimental observation of snapping frequency, $\omega$, as a function of non-reciprocity, $k^{\mathrm{o}}$, Fig. 3(c).
Figure 4: Polyfunctions of non-reciprocal filaments. The same filament exhibits different locomotion modalities in response to distinct environmental perturbations. (a, b) Buckling our filaments via a free-standing brace, unidirectional shape cycles enable crawling on a substrate. Key frames are shown in panel (a) (i-v), with projections onto the lowest shape modes in (b). First, the rear linkage descends, driving the chain forward (i, ii). The filament’s anchor point then swaps to the front linkage (iii), as the rear retracts (iv) and the cycle repeats (v). Displacement measured in body lengths (BL). (c) Locomotion velocity and snapping frequency versus the non-reciprocity $k^{o}$. The filament switches gait from crawling to jumping as the snapping frequency reaches the timescale of vertical motion in the filament’s centre of mass. (d) The active filament navigates an obstacle by spontaneously switching gaits. The trajectory of the centre of mass is shown here and colored by the snapping frequency. As the filament touches the obstacle and leaves the substrate, it switches to the jumping gait. Once the middle units touch the obstacle, the filament snaps slowly again and lifts its body to climb up the obstacle. Inset shows details of the transition where the filament jumps, climbs, then jumps over the obstacle. (e) A half-clamped filament spontaneously steps back and forth when driven into contact with a substrate. Inset shows a kymograph with the angle deviations at free vertices over time, which clearly shows unidirectional waves. (f) A standing active filament digs into a granular pile of steel beads. While compressing the active filament, it scoops steel beads and pulls them aside as it snaps. (g) Breaking left–right symmetry enables the standing filament to walk. We tilt the filament by an angle $\phi$ and allow the top end to slide freely. The filament walks persistently towards the tilting direction. (h) Walking velocity versus the tilt angle $\phi$.
We have seen that buckling filaments at sufficiently high $k^{\mathrm{o}}$ drives them into a limit cycle of shape changes. We now use this mechanistic insight to enable multiple modalities of locomotion in a single filament. We first buckle our chain with a brace that constrains its length while allowing it to translate freely in space [Fig. 4(a), Supplementary Video 4]. Placed upon a substrate, we find that the shape cycles executed by the beam now translate to a crawling state of locomotion [Fig. 4(b)]. The direction of motion is aligned with the sense of wave amplification: The beam crawls via a wave that originates from its rear and propagates to its head, analogously to the metachronal waves of the caterpillar [38] or the lagged phase of microswimmers [39, 40].
Our crawler is powered by an external brace, which maintains the filament in a buckled state. We now allow the environment itself to cause buckling, driving a feedback loop between filament snapping and environmental perturbation that will autonomously select a locomotion gait. We clamp a filament at one end, and drive the unconstrained end into a substrate. The filament buckles and begins to snap. However, because the open end is free to leave the substrate, the entire chain now steps back and forth [Fig. 4(e), Supplementary Video 4].
Driving our free chain into a granular pile of steel beads, we find that chain snapping digs a hole in the pile [Fig. 4(f), SM §VI]. Excavation ends when the filament has dug enough space to return to its rest length. Because we have broken the symmetry of the boundary conditions applied to the chain, we also induce an additional polarity running from the clamped end to the free end. This polarity breaks the symmetry $k^{\mathrm{o}}\rightarrow-k^{\mathrm{o}}$: A wave amplifying from the clamped end to the free end results in a different gait to a wave amplifying from free to clamped. When waves run in to the granular pile, we find a burrowing gait and narrower hole. By contrast, when the wave begins within the pile, we find larger sweeping motions and a wider excavation [Supplementary Video 4].
The interplay between buckling via an external load and unidirectional beam snapping suggests that our active filaments would make natural walkers. However, despite breaking $k^{\mathrm{o}}\rightarrow-k^{\mathrm{o}}$ symmetry along the beam, the geometry shown in Fig. 4(d) preserves a left/right reflection symmetry perpendicular to the beam, prohibiting net translation along this axis—the walker takes a step forward, but an equal step back. However, upon tilting the filament relative to the substrate by an angle $\phi$ we break this symmetry and immediately find an asymmetric walking gait and net translation [Fig. 4(g), Supplementary Video 4], with a walking velocity which scales as $\phi$ [Fig. 4(h)]. Analogously to the digger, the gait of the walker depends on sense of wave propagation, and we find that only waves leading from the foot result a persistent walking gait.
We have demonstrated that buckling in materials which break reciprocity enables a new class of instabilities: not only snapping between static states, but also self-sustained oscillations. These oscillations occur in free-standing structures, in contrast to previous work on active filaments [25, 26, 27, 28, 29], a crucial feature that we exploit to engineer self-supported locomotors which crawl, dig, and walk. Passive beam buckling has underpinned a generation of structures and architected materials that achieve functionality through nonlinearity [42, 43, 44, 45]. We anticipate that active buckling represents a similarly fruitful pathway towards unconventional soft robots and multistable active matter.

### v2

@@ I The continuum mechanics of a non-reciprocal beam
where $\lambda$ is the growth rate of the spatial Fourier mode with wavenumber $q$.Equation (23) appears as (3) of the Main Text.
II Analysis of the odd von Mises truss
II.1 Derivation of the equations of motion via the principle of virtual work
The odd active torque for each of the vertices is given by
$\tau_{\text{L}}$ $=k^{\text{o}}(\theta_{\text{LC}}-\pi)\text{,}$ (13)
$\tau_{\text{R}}$ $=-k^{\text{o}}(\theta_{\text{RC}}-\pi)\text{,}$ (16)
$E_{s}=$ $\frac{1}{2}k\bigg(l_{\text{0}}-\big[(-\text{d}l+l-w_{\text{L}}-w_{\text{R}})^{2}$
$\qquad+(w_{\text{L}}\tan(\theta_{\text{L}})-w_{\text{R}}\tan(\theta_{\text{R}}))^{2}\big]^{1/2}\bigg)^{2}$
$\mathcal{R}=\frac{\Gamma}{2}\left[\dot{\theta}_{\text{L}}^{2}+\dot{\theta}_{\text{LC}}^{2}+\dot{\theta}_{\text{RC}}^{2}+\dot{\theta}_{\text{R}}^{2}\right]+$ (19)
$k^{\text{o}}[(\theta_{\text{LC}}-\pi)\dot{\theta}_{\text{L}}+(\theta_{\text{RC}}-\pi-\theta_{\text{L}})\dot{\theta}_{\text{LC}}$
$+$ $(\theta_{\text{R}}-\theta_{\text{LC}}+\pi)\dot{\theta}_{\text{RC}}-(\theta_{\text{RC}}-\pi)\dot{\theta}_{\text{R}}]\text{.}$
Here $\Gamma$ is the hinge dissipation and $k^{\text{o}}$ the microscopic non-reciprocity. Terms like $\frac{\Gamma}{2}\dot{\theta_{L}}^{2}$ in (45) are the viscous power dissipated by hinge $\theta_{L}$. Terms like $k^{\mathrm{o}}(\theta_{\text{LC}}-\pi)\dot{\theta}_{\text{L}}$ are the power generated by the odd torques acting on vertex $\theta_{\text{L}}$. Given a Lagrangian $L(\theta,\dot{\theta})$ that encodes conservative dynamics, and a dissipation function $\mathcal{R}(\theta,\dot{\theta})$, the Rayleighan formalism gives equation of motion via
In the truss, the Lagrangian $L=-(E_{s}+E_{b})$ where the potential energy is the sum of bending and stretching energies (17), (18), neglecting inertia. The dissipation $\mathcal{R}$ is given by (45).
$\theta_{\text{LC}}$ $=\pi-\theta_{\text{L}}-\text{arctan}\left(-\text{d}l+l-w_{\text{L}}-w_{\text{R}},w_{\text{L}}\tan(\theta_{\text{L}})-w_{\text{R}}\tan(\theta_{\text{R}})\right)\text{,}$
$\theta_{\text{RC}}=$ $\pi-\theta_{\text{R}}+\text{arctan}\left(-\text{d}l+l-w_{\text{L}}-w_{\text{R}},w_{\text{L}}\tan(\theta_{\text{L}})-w_{\text{R}}\tan(\theta_{\text{R}})\right)\text{.}$
We also non-dimensionalize energies by $B$ and lengths by $l=3l_{0}$. Time will be renormalized by $\Gamma/B$. Thus our dimensionless parameters are given by
$k^{\text{o}}\leftarrow\frac{k^{\text{o}}}{B}\text{,}$ (25)
$k\leftarrow\frac{kl^{2}}{B}\text{,}$ (26)
$\text{d}l\leftarrow\frac{\text{d}l}{l}\text{.}$ (27)
Expanding to $4^{\text{th}}$ order in $\epsilon$ gives the following energy:
$+\frac{1}{576}\epsilon^{4}\Big(144\text{d}l^{2}k-2592\text{d}l\theta_{A}^{2}-24\text{d}l\theta_{A}^{2}k-96\text{d}l\theta_{S}^{2}k$
$\qquad-288\text{d}lkw_{\text{S}}-432\theta_{A}^{4}+1728\theta_{A}^{2}\theta_{S}^{2}+9\theta_{A}^{4}k$
$\qquad+24\theta_{A}^{2}\theta_{S}^{2}k+16\theta_{S}^{4}k+144kw_{\text{A}}^{2}-96\theta_{A}\theta_{S}kw_{\text{A}}$
$\qquad+432kw_{\text{S}}^{2}-72\theta_{A}^{2}kw_{\text{S}}+96\theta_{S}^{2}kw_{\text{S}}-5184\theta_{A}\theta_{S}w_{\text{A}}$
$\qquad+7776\theta_{A}^{2}w_{\text{S}}\Big)\text{,}$ (28)
and the dissipation is given by
$\mathcal{R}=\frac{1}{2}\epsilon^{2}\left(10k^{\text{o}}\theta_{\text{S}}\dot{\theta}_{\text{A}}-10k^{\text{o}}\theta_{\text{A}}\dot{\theta}_{\text{S}}+5\dot{\theta}_{\text{A}}^{2}+4\dot{\theta}_{\text{S}}^{2}\right)$
$\qquad+\frac{1}{4}\epsilon^{4}\Big(-72\text{d}lk^{\text{o}}\theta_{\text{S}}\dot{\theta}_{\text{A}}+72\text{d}lk^{\text{o}}\theta_{\text{A}}\dot{\theta}_{\text{S}}-54\text{d}l\dot{\theta}_{\text{A}}^{2}$
$\qquad-12k^{\text{o}}\theta_{\text{A}}^{2}\theta_{\text{S}}\dot{\theta}_{\text{A}}+16k^{\text{o}}\theta_{\text{S}}^{3}\dot{\theta}_{\text{A}}+4k^{\text{o}}\theta_{\text{A}}^{3}\dot{\theta}_{\text{S}}+16k^{\text{o}}\theta_{\text{A}}\theta_{\text{S}}^{2}\dot{\theta}_{\text{S}}$
$\qquad-48k^{\text{o}}\theta_{\text{S}}^{2}\dot{w}_{\text{A}}+72k^{\text{o}}\theta_{\text{A}}\theta_{\text{S}}\dot{w}_{\text{S}}+72k^{\text{o}}\theta_{\text{S}}w_{\text{S}}\dot{\theta}_{\text{A}}$
$\qquad-72k^{\text{o}}\theta_{\text{A}}w_{\text{S}}\dot{\theta}_{\text{S}}+24\theta_{\text{A}}\theta_{\text{S}}\dot{\theta}_{\text{A}}\dot{\theta}_{\text{S}}+12\theta_{\text{S}}^{2}\dot{\theta}_{\text{A}}^{2}-9\theta_{\text{A}}^{2}\dot{\theta}_{\text{A}}^{2}$
$\qquad-36\theta_{\text{S}}\dot{\theta}_{\text{A}}\dot{w}_{\text{A}}-36w_{\text{A}}\dot{\theta}_{\text{A}}\dot{\theta}_{\text{S}}+54\theta_{\text{A}}\dot{\theta}_{\text{A}}\dot{w}_{\text{S}}+54w_{\text{S}}\dot{\theta}_{\text{A}}^{2}\Big)\text{.}$ (29)
$\frac{\partial}{\partial w_{\text{{S,A}}}}(E_{b}+E_{s})=0,$ (30)
which gives
$w_{\text{A}}=\frac{(k+54)\theta_{\text{A}}\theta_{\text{S}}}{3k},$ (31)
We then set $\dot{w}_{\text{S,A}}=0$ in the dissipation functional (29) and calculate the dynamical equations for only the two angular degrees of freedom. These equations are found by application of the variational principle (20) as
which to third order in $\epsilon$ gives
Equation (34) is of the general cubic form given in (7) of the Main Text, namely
@@ II.1 Derivation of the equations of motion via the principle of virtual work
As described in the Main Text, we note that the $\mathbb{Z}_{2}$ symmetry $(\theta_{S},\theta_{A})\rightarrow(-\theta_{S},-\theta_{A})$ forces nonlinear terms of cubic order only in (27). The symmetry under simultaneous reflection and a reversal in the sense of non-reciprocity, $(\theta_{S},\theta_{A},k^{\mathrm{o}})\rightarrow(\theta_{S},-\theta_{A},-k^{\mathrm{o}})$ constrains $k^{\mathrm{o}}$ to appear only in the off-diagonal elements of $J$ in (28), and enforces the alternating pattern of terms proportional to $k^{\mathrm{o}}$ in ${\bf A}_{i}$. More formally, we split these terms into ${\bf A}_{i}={\bf A}_{i}^{\mathrm{passive}}(k)+{\bf A}_{i}^{\mathrm{odd}}(k^{\mathrm{o}},k)$. We require the symmetry ${\bf A}^{\mathrm{odd}}_{i}=(-1)^{i}\sigma_{z}{\bf A}^{\mathrm{odd}}_{i}$, with $\sigma_{z}$ a Pauli matrix. We now analyze the bifurcation structure of the dynamics  (27).
II.2 Bifurcation structure of the odd von Mises truss
Figures 3(f, g) show a selection of phase portraits for the truss dynamics (27) as the non-reciprocity $k^{\mathrm{o}}$ and compression $\text{d}l$ are varied. In the limit of no active force, $k^{\text{o}}=0$, we find classic symmetrically buckled solutions
$\theta_{\text{S}}^{\star}=\pm\frac{\sqrt{3}\sqrt{\text{d}lk-18}}{\sqrt{k}},\quad\theta_{\text{A}}^{\star}=0\text{.}$ (37)
As $k^{\mathrm{o}}$ is increased, these buckled fixed points merge in a SNIC bifurcation, and a limit cycle is born.
To locate the critical non-reciprocity $k_{c}^{o}$ and compression $\text{d}l_{c}$ for the CEP shown in Fig. 3 of the Main Text, we demand a double zero in the eigenvalues of the Jacobian $J$ in (28). We enforce this condition by solving for $\text{Tr}({J})=\text{Det}({J})=0$, which gives
$k^{\text{o}}_{\text{c}}=\pm\frac{k+162}{\sqrt{20k^{2}-2160k+25515}},$ (38)
$\text{d}l_{\text{c}}=\frac{45}{2k-81}\text{.}$ (39)
$\text{d}l_{\text{c}}=\frac{45}{2k}\text{,}$ (41)
which is the location of the CEP given in the Main Text. Substituting these critical values back into the Jacobian (28), perturbing as $k^{\text{o}}=k^{\text{o}}_{\text{c}}+\delta k^{\text{o}}$, $\text{d}l=\text{d}l_{\text{c}}+\delta l/k$ and taking the limit $k\to\infty$ we find the normal form
which is given as (8) of the Main Text. In order to find the theoretical snapping frequencies shown in Fig. 3(e) of the Main Text, we look for the imaginary part of eigenvalues, $\lambda_{J}$ of $J$ in the vicinity of the CEP:
III Construction of our robotic metamaterials
$\tau^{a}_{i}=\begin{cases}k^{\mathrm{o}}\left(\delta\theta_{i+1}-\delta\theta_{i-1}\right)\mathrm{,}|\delta\theta_{i+1}-\delta\theta_{i-1}|<\tau_{\mathrm{max}}/k^{\mathrm{o}},\\ \mathrm{sgn}(\delta\theta_{i+1}-\delta\theta_{i-1})\tau_{\mathrm{max}}\mathrm{,}|\delta\theta_{i+1}-\delta\theta_{i-1}|>\tau_{\mathrm{max}}/k^{\mathrm{o}},\end{cases}$ (44)
[…]
The torques and angles in (44) are implemented using arbitrary digital units—bits—that are specific to our robotic implementation. We calibrate these digital values against standard physical units by measuring torque-displacement slopes for different values of the electronic feedback, as described in Ref. [53]. We find a conversion factor $846\text{\,}\mathrm{mN}\text{\,}\mathrm{m}\text{\,}{\mathrm{rad}}^{-1}\text{\,}{\mathrm{bit}}^{-1}$ taking us from digital to physical non-reciprocities $k^{\mathrm{o}}$. The coreless motor saturates at a maximum torque of $\tau_{\mathrm{max}}=$12\text{\,}\mathrm{mN}\text{\,}\mathrm{m}$\$. Each link has a length of $l_{0}=7.5\,\mathrm{cm}$. Adjacent vertices are also connected by viscoelastic hinges, whose materials properties we vary depending on the experimental protocol (§IV). For further details, see Ref. [53].
Figure 5: Response Time Calibration. We estimate the timescale for the process of digital data acquisition, signaling, and motor actuation by applying an impulse to a 2-bar linkage, filmed under a high speed camera. We find a conservative upper bound of $50\text{\,}\mathrm{ms}$, which is $40\times$ faster than the most rapid oscillations our beams.
IV Experimental protocol
Figure 6: Details of our experimental buckling setup. A typical setup during buckling experiments on our robotic metamaterials, showing a chain of 12 units clamped to an air table at either end, with an Agilus 30 viscoelastic polymer scaffold attached to the top of the chain to provide hinge elasticity and dissipation. Scale bar $10\text{\,}\mathrm{cm}$.
All experiments take place on top of a custom-made low-friction air table to minimize stick-slip substrate friction. The table consists of two $1.5\times 1.5$m Plexiglass plates that sandwich 5-mm-wide air channels. The top plate is pierced by an array of holes (pitch 10 mm, diameter 1 mm) conveying air pressurized at 10 bar. Each motorized vertex floats on a thin layer of pressurized air without making contact with the table. The experiments are filmed using a Nikon D5600 camera equipped with a 50-mm lens, recording 2-Mpx images at 50 frames per second. A marker is placed on the top of each vertex and we use the Python package OpenCV to detect and track the position of each vertex with a spatial resolution of 0.8 mm. We show a typical table setup in Fig. 6.
To provide torsional rigidity and viscous dissipation to the hinges of our robotic metamaterials, we use two different viscoelastic polymers. For the data shown in Fig. 3, we use a silicone rubber with a thickness of $2\text{\,}\mathrm{mm}$. This rubber hinge is well described with a Kelvin-Voigt model, with passive elastic torsional stiffness of $B=$20\text{\,}\mathrm{mN}\text{\,}\mathrm{m}\text{\,}{\mathrm{rad}}^{-1}$\$ and viscous dissipation $\Gamma=$0.4(1)\text{\,}\mathrm{mN}\text{\,}\mathrm{m}\text{\,}\mathrm{s}\text{\,}{\mathrm{rad}}^{-1}$\$. See Ref. [53] for calibration details. The simple viscoelastic response of this rubber means that it is well suited for the detailed calibration performed in Fig. 3. However, for longer chains (Figs. 14) we find that the dissipation of this rubber is insufficient to drive us into the viscous regime in which snapping is dominated only by the lowest modes of a longer chain.
For our chain snapping (Fig. 1) and locomotion (Fig. 4) experiments, we therefore use a highly viscous polymer. We 3D print hinges from a viscoelastic rubber-like PolyJet photopolymer (Stratasys Agilus 30). The viscoelastic response of this material has been characterised via a generalized Maxwell model in, for example, Ref. [16]. Complex viscoelastic response makes the Agilus 30 ill-suited for the detailed calibration shown in Fig. 3. However, its increased viscous dissipation renders it ideal for our locomotion experiments. In this study, we used two thicknesses of Agilus 30 bands: 3 mm and 6 mm. We determined an approximate angular stiffness $B$ and angular dissipation $\Gamma$ by placing a single robotic vertex in a rotational testing device (Instron E3000 linear-torsional) equipped with a torsional load cell (angular resolution: 2.5 arcsecs; torque resolution: 1 mN m). To determine the angular stiffness $B$ of the rubber band, the torque was measured as a function of angular deviation. The value of $B$ is taken to be the slope of the linear fit. We find $B\approx 3.90\ \mathrm{mN\cdot m/rad}$ and $B\approx\ 8.78\ \mathrm{mN\cdot m/rad}$ for the rubber band with a thickness of 3 mm and 6 mm, respectively (Fig. 7). To determine angular dissipation $\Gamma$, we place the rubber band on the linkage system, and we apply an angular perturbation to the free edge. We measure the angular displacement and extract $\Gamma$ from the decay rate. We find $\Gamma\approx 7.38\ \mathrm{mN\cdot m\cdot s/rad}$ and $\Gamma\approx 22.04\ \mathrm{mN\cdot m\cdot s/rad}$ for the rubber band with a thickness of 3 mm and 6 mm, respectively (Fig. 8). We emphasise that these measurements use a Kelvin-Voigt model for the Agilus 30, which is an oversimplification [16]. However, they indeed demonstrate the dramatically increased dissipation $\Gamma$ of the Agilus relative to a silicone rubber.
Figure 7: Calibration of angular stiffness $B$ for our Agilus-30 hinges. The restoring torque of the building block with a viscous Agilus band measured over a range of twist angles $\delta\theta$. The slope of the linear regression yields a value $B\approx 3.90\ \mathrm{mN\cdot m/rad}$ and $B\approx\ 8.78\ \mathrm{mN\cdot m/rad}$ for the rubber band with a thickness of (a) 3 mm and (b) 6 mm.
Figure 8: Calibration of angular damping coefficient $\Gamma$ for our Agilus-30 hinges. The decaying oscillations of the building block with a viscous Agilus band were fitted to the solution of a damped harmonic oscillator. We find viscous damping coefficients of $\Gamma\approx 7.38\ \mathrm{mN\cdot m\cdot s/rad}$ and $\Gamma\approx 22.04\ \mathrm{mN\cdot m\cdot s/rad}$ for the Agilus rubber band with a thickness of (a) 3 mm and (b) 6 mm.
IV.1 Unbuckled chain
We assemble a 22-vertex chain, damped using $6\text{\,}\mathrm{mm}$ Agilus-30 hinges, loosely confined at either end with masking tape, and not under compression or tension. The non-reciprocity $k^{\mathrm{o}}=34\mathrm{mN\cdot m\cdot s/rad}$, and a manual perturbation is applied to the chain’s edge, and midpoint [Fig. 2(a, b)].
IV.2 Buckled chain
For the chain buckling experiments shown in Figs. 16, we assemble an open 12-vertex filament, with each hinge damped using the $6\text{\,}\mathrm{mm}$ Agilus 30. The boundary conditions mimic a clamped-tangent boundary condition in classical buckling: At each end of the chain, the two extremal motors ($i=1,2,11,12$) are clamped in position. The boundary motors $(i=1,12)$ are clamped to external rigid bars via a custom 3D-printed pinch, which is displaced and position controlled to apply the external compression $\text{d}l$. The boundary conditions on the non-reciprocal torques mimic those implemented in our discrete theory §II. For $i=1,12$, the active torque $\tau_{i}=0$. For $i=2$, $\tau_{2}=k^{\mathrm{o}}\delta\theta_{3}$. For $i=11$, $\tau_{11}=-k^{\mathrm{o}}\delta\theta_{10}$.
IV.3 Buckled von Mises truss
[…]
IV.4 Crawling experiments
Figure 9: Details of crawling experiment setup. A chain of 6 units confined via a laser cutting PMMA brace. By changing the distance between the edge unit and fastening the screws, we can maintain the chain in a buckled state. An Agilus 30 viscoelastic polymer scaffold is attached to the top of the chain to provide hinge elasticity and dissipation. Additional pieces of silicone rubber are put on the side of each unit to ensure frictional contact with the substrate. Scale bar $10\text{\,}\mathrm{cm}$.
Figure 10: Details of walking experiment setup. A chain of 6 units whose the top-end unit is clamped to a slide rail and the bottom-end unit can freely move. The chain is tilted to an angle $\phi$ by fastening the screw on the orange brace. An Agilus 30 viscoelastic polymer scaffold is attached to the top of the chain to provide hinge elasticity and dissipation. Scale bar $10\text{\,}\mathrm{cm}$.
Figure 11: Crawling and stepping are robust to perturbations. We manually perturb (a) our crawling chain and (b) the stepping motion that underpins walking and digging. In the face of repeated perturbations from different angles we find that each dynamics is attracted back to the same limit cycle (see also Supplementary Video 5). (a)(i) and (b)(ii) show selected snapshots of this perturbation and recovery process. Symmetric and antisymmetric modes defined as in Fig. 4.
IV.5 Digging experiments
IV.6 Walking experiments
IV.7 Additional Robustness Tests
To further check the robustness of crawling, digging and walking (Fig. 4), we manually perturb both the limit cycle of our crawler, and the limit cycle for the fundamental stepping motion which underpins walking and digging (Fig. 11, Supplementary Video 5). We observe that the dynamics is repeatedly attracted back to the same limit cycle in the face of multiple perturbations.